OpenStudy (anonymous):

Hookes Law. The distance d when a spring is stretched by a hanging object varies directly as the weight w of the object. If the distance is 32 cm when the weight is 4 kg, what is the distance when the weight is 9 kg?

OpenStudy (anonymous):

d=kw where k is a constant. 32=k(4) Find k. 9=k(?)

OpenStudy (a_clan):

32=k(4) k=32/4 = 8 cm/kg d=k(9) d=8*(9) = 72 cm